OpenStudy (anonymous):

Question 3. 3. What is 40% of 260? (Points : 1) 15 104 156 364

jagr2713 (jagr2713):

its C

jagr2713 (jagr2713):

100-40=60 260*0.60=156

OpenStudy (anonymous):

@jagr2713 Thats 60 % not 40%.

OpenStudy (anonymous):

(260/100 )*40 = 104

jagr2713 (jagr2713):

Oops sorry i didn't read it correctly. I apologize

OpenStudy (anonymous):

 104